OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 2x + 3y = 18?

OpenStudy (anonymous):

mathhow.com or something like that. Google it, trust me.

hartnn (hartnn):

try to isolate y like bring it in the form y=mx+c

hartnn (hartnn):

for that, first subtract 2x from both sides, what u get ?

OpenStudy (anonymous):

3y=18-2x

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes, now to get y by itself, just divide 3 on both sides, what u get ?

OpenStudy (anonymous):

y=6-2x

hartnn (hartnn):

you for got to divide 3 to the term 2x you should have got -2x/3 instead of just -2x

OpenStudy (anonymous):

ohh ok but how do i divide 2x by 3

hartnn (hartnn):

don't! keep it as it is ... so, you have y = -2/3 x +6 comparing with y=mx+c you get slope = m = -2/3 and y-intercept = c = 6 got it ?

OpenStudy (anonymous):

Yea